English

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(not A).

Advertisements
Advertisements

Question

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(not A).

Sum
Advertisements

Solution

P(A) = 0.42, P(B) = 0.48

P(A and B) = P(A ∩ B) = 0.16

P(not A) = P(A’) = 1 – P(A)

= 1 – 0.42

= 0.58

shaalaa.com
Concept of Probability - Probability of 'Not', 'And' and 'Or' Events
  Is there an error in this question or solution?
Chapter 14: Probability - EXERCISE 14.2 [Page 307]

APPEARS IN

NCERT Mathematics [English] Class 11
Chapter 14 Probability
EXERCISE 14.2 | Q 17. (i) | Page 307

RELATED QUESTIONS

If E and F are events such that P(E) = `1/4`, P(F) = `1/2` and P(E and F) = `1/8`, find

  1. P(E or F)
  2. P(not E and not F).

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P (not B)


In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.


In a class of 60 students, 30 opted for NCC, 32 opted for NSS and 24 opted for both NCC and NSS. If one of these students is selected at random, find the probability that

  1. The student opted for NCC or NSS.
  2. The student has opted neither NCC nor NSS.
  3. The student has opted NSS but not NCC.

Out of 100 students, two sections of 40 and 60 are formed. If you and your friend are among the 100 students, what is the probability that

  1. you both enter the same sections?
  2. you both enter the different sections?

Three letters are dictated to three persons and an envelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.


A dice is thrown. Find the probability of getting a prime number


A dice is thrown. Find the probability of getting a multiple of 2 or 3.

 

In a simultaneous throw of a pair of dice, find the probability of getting:

8 as the sum


In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers


In a simultaneous throw of a pair of dice, find the probability of getting an even number on one and a multiple of 3 on the other


In a simultaneous throw of a pair of dice, find the probability of getting neither 9 nor 11 as the sum of the numbers on the faces


In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 6


In a simultaneous throw of a pair of dice, find the probability of getting a sum less than 7


In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 7


In a simultaneous throw of a pair of dice, find the probability of getting a total of 9 or 11


In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.

 

Three coins are tossed together. Find the probability of getting exactly two heads


Three coins are tossed together. Find the probability of getting at least two heads


Two dice are thrown. Find the odds in favour of getting the sum 5.

 

 


A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white and odd numbered .


Fill in the blank in the table:

P (A) P (B) P (A ∩ B) P(A∪ B)
\[\frac{1}{3}\] \[\frac{1}{5}\] \[\frac{1}{15}\] ......

If A and B are two events associated with a random experiment such that
P(A) = 0.5, P(B) = 0.3 and P (A ∩ B) = 0.2, find P (A ∪ B).


If A and B are two events associated with a random experiment such that
P (A ∪ B) = 0.8, P (A ∩ B) = 0.3 and P \[(\bar{A} )\]= 0.5, find P(B).

 


There are three events A, B, C one of which must and only one can happen, the odds are 8 to 3 against A, 5 to 2 against B, find the odds against C


One of the two events must happen. Given that the chance of one is two-third of the other, find the odds in favour of the other.


In a single throw of two dice, find the probability that neither a doublet nor a total of 9 will appear.


A person write 4 letters and addresses 4 envelopes. If the letters are placed in the envelopes at random, then the probability that all letters are not placed in the right envelopes, is


If the probability of A to fail in an examination is \[\frac{1}{5}\]  and that of B is \[\frac{3}{10}\] . Then, the probability that either A or B fails is

 
 

Three integers are chosen at random from the first 20 integers. The probability that their product is even is


Five persons entered the lift cabin on the ground floor of an 8 floor house. Suppose that each of them independently and with equal probability can leave the cabin at any floor beginning with the first, then the probability of all 5 persons leaving at different floor is


A box contains 10 good articles and 6 with defects. One item is drawn at random. The probability that it is either good or has a defect is


In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy two tickets.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×